PREPARING YOUR CAMERA-READY ARTICLE

FOR OFFSET REPRODUCTION

John Q. Editor,^{1} Jane Doe,^{2} and Richard Smith^{1}

1 Accelerator Design and Operations Division 2 Magnet Division Superconducting Super Collider Laboratory ^{∗} Dallas, Texas 75237

INTRODUCTION

The purpose of this leaflet is twofold: (1) to furnish an example showing just how the various elements of a technical paper should be presented, and (2) to provide a brief summary of the rules governing the preparation of camera-ready copy. It is of course not possible to address all the problems that may arise in these few pages. If you run into a problem that you cannot solve, call us at (212) 620-8042. Fax (212) 463-0742.

We must distinguish between mechanical requirements for the preparation of copy suitable for direct offset reproduction and editorial guidelines concerning such matters as style, content, and scope. Editorial guidelines are the responsibility of the volume editor, and questions about such matters should be addressed to him. Mechanical requirements are the province of the publisher and constitute the principal subject of these guidelines.

PREPARING THE CAMERA-READY ARTICLE

Acceptable copy can be prepared using various computer-based systems or even an ordinary electric typewriteer. However, the guidelines that follow were written with the assumption that your article will be prepared using a computer-based system incorporating a printer capable of printing proportionally-spaced serif (booklike) type. A book-like typeface should be used if available, but articles prepared in a monospaced (typewriter-style) typeface are acceptable. If your article will be printed in a monospaced typeface, please refer to our supplement, “If You Are Using a Typewriter.”

∗ Operated by the Universities Research Association, Inc., for the U.S. Department of Energy under

Contract No. DE-AC35-89ER40486.

Please note that the output device used to generate the article must be a laser printer, letter-quality dot-matrix printer, or phototypesetter. Draft and near-letter-quality printers do not produce type of acceptable quality.

The trim size of the printed volume will by 6-1/2 × 9-7/8 inches (ca. 165 × 251 mm), with a final type area—the area encompassing the text, figures, tables, etc.—of 30-1/2 × 50-1/2 picas (ca. 5-1/16 × 8-3/8 inches or 128 × 213 mm). In order to improve the type quality in the printed volume, the camera copy you will be preparing will be designed to be reduced to 82-1/2% of its initial size during the printing process. Thus, the type area of your camera-ready article should be 37 × 61 picas (ca. 6-1/8 × 10-1/8 inches or 156 × 257 mm).

The article may be printed on any good-quality white paper, but if possible use legal-size paper, which will allow margins of greater than one inch to be left around the type area.

TYPE STYLE AND TEXT

The article should be printed in 12-point Times Roman. If Times Roman is not available, choose a typeface similar to Times Roman, e.g., Times or Dutch. If no such typeface is available, the article should be printed in another of the common roman, serif typefaces, such as the ones listed in the first column of Table 1. If 12-point type is not available, 10- or 11-point type is also acceptable. Sans serif typefaces such as Helvetica and Univers should not be employed as main text faces, though sans serif fonts are excellent choices for labeling figures.

Table 1. Examples of acceptable and unacceptable alternatives should no typeface similar to Times Roman be available.

Acceptable  Unacceptable
Proportionally-spaced^{1} Monospaced^{2} San serif  Special^{3}
Bookman  Courier  Avante Garde Boldface fonts
Palatino  Helvetica  Italic fonts
Schoolbook  Script

1 The main text typeface should be no greater than 12-point and no smaller than 10-point. 2 For special instructions regarding articles prepared in monospaced (typewriterstyle) typefaces, see “If You Are Using a Typewriter.” 3 Boldface and italic type may, of course, be used in specialized applications, as in headings and where emphasis is desired.

In most cases, if 12-point type is used, the space between lines hould be about two points: the default line spacing specified by most word processing packages is satisfactory. The line spacing should be constant throughout the article. To prevent the line spacing from varying, it may be necessary to set a fixed line spacing instead of letting your word processing package adjust the line spacing automatically. If the text will contain many subscripts, superscripts, or oversize symbols, the article line spacing should be increased slightly, but double-spaced text is not acceptable.

The text must be clearly printed and uniformly dark, without noticeable streaks or fading. Please check to make sure that the text near the margins is as dark as the text near the middle of the page and that the spaces within the small characters are not filled with ink.

The right margin of the text should be justified (made even). However, the right margins of display elements, such as the text subheadings, article title, author names, author addresses, figure captions, table titles, table columns, and references should be left ragged. Flaws in the justification frequently appear when minor revisions are made to the article.

The first line of each paragraph should be indented 2 picas (ca. 1/3 inch or 8.5 mm). No extra space should be left between paragraphs.

Maintaining the proper text-area dimensions is important. The text should be run the full width of the type area as described above, and each page should be as close to the full text-area length as possible. Articles in which the text is much narrower or wider than 37 picas or in which many of the pages are short of full length without an evident reason (such as a large equation or a text subheading at the top of the succeeding page) are not acceptable (Figure 1).

TEXT SUBHEADINGS

In most cases, only two values of subheadings should be employed. The paragraphs below demonstrate the use of such subheadings. All subheading should be set in type fo the same typeface and size as the text.

FIRST VALUE SUBHEADINGS

These headings should be printed in bold capital letters, on a separate line, flush with the left margin. Leave a 2-line space above and a 1-line space below first-value subheadings.

Second-Value Subheadings

These headings should be printed in bold upper- and lower-case type, with the initial letter of each major word capitalized, on a separate line, flush left. Leave f1-line spaces above and below second-value subheadings.

Third-Value Subheadings. Third-value subheadings, if unavoidable, should follow the second-value subheadings in type style but should be run into the text. A paragraph beginning with a third-value subheading should be preceded by a 1-line space.

REFERENCES

The list of works cited should appear at the end of the article. Establishing rules for styling the references is the province of the editor, so questions concerning reference styling should be addressed to the editor. If the editor has not chosen to establish style guidelines for the references, please follow the examples below:

References by Name and Year

If reference citation is by name and year, the text citation may take one of the following forms: “… as shown by Miller (1967), the … ” or “… has often been demonstrated (Smith and Jones, 1972; Brown et al., 1974) that ….” In this case, the reference list at the end of the paper must be in alphabetical order of the first authors’ names and presented in the following style:

Brown, C.D., Green, M.P., and Robinson, S.A., 1974, Article title with only the first word having

an initial capital, Journal Name Abbr. 37:468. Miller, R.J., 1967, “Book Title with the Initial Letter of Each Major Word Capitalized,” Pub-

lisher, City. Smith, A.B., and Jones, C.D., 1972, Article or chapter title, in: “Book Title,” W.F. White, ed.,

Publisher, City.

Numbered References

If reference citation is by number, the text citation may take one of the following forms: “… as shown by Miller,^{1} the …,” “… has often been demonstrated^{2,3} that ….” In this case, the references are to be numbered in the order in which they are cited in the text and presented int eh following style:

1. R.J. Miller. “Book Title with the Initial Letter of Each Major Word Capitalized,” Publisher, City (1967).

2. A.B. Smith and C.D. Jones, Article or chapter title with only the first work having an initial capital, in: “Book Title,” W.F. White, ed., Publisher, City (1972).

3. C.D. Jones, M.P. Green, and S.A. Robinson, Article title, Journal Name Abbr. 37:468 (1974).

MISCELLANEOUS

The Opening Page

The Opening page should follow the example shown in these guidelines. There should be a ten-line space above the article title, and the title should be printed in capital letters. In most cases, lines in the article title should not extend more than two-thirds page width. The author names and addresses should be printed in upperand lower-case type and should be indented from the left margin about 1 inch or 250 mm. The text or first subheading should start three lines below the addresses, not on the following page.

Equations

Equations should either be centered in the page width or indented 2 picas from the left margin. Leave 1-line spaces above, below, and between equations. Equations numbers hould be placed in parentheses, flush with the right margin. Equations should be typed or printed—handwritten equations are not acceptable. For example:

Separating variables and expanding the current density, we obtain

\frac{I^{2}(t)}{s^{2}}dt = μ\frac{C(T)}{ρ(T)}dT  (1)

where s = cross sectional surface area of the rib cage.

Rearranging to solve for s^{2} and integrating, we obtain

s^{2} = \frac{∫_{0}^{∞} I^{2}(t)dt}{μ ∫_{T_{init}}^{T_{max}} \frac{c(T)}{ρ(T)}dT} .  (2)

Footnotes and Endnotes

Notes may appear at the bottom of the pertinent pages or gathered together in a notes section immediately preceding the reference section. ^{†}

Acknowledgements

Acknowledgements, if required, should appear in a section immediately before the reference section, or the endnote section, if there is one.

Special Symbols

Special symbols that cannot be printed may be hand-drawn using a fine black pen. Symbols drawn with pens of other colors, with pencils, or with felt tip pens will not reproduce well. Symbols should be drawn roughly the same size as the printed characters surrounding them, and the spacing around symbols should be similar to that around ordinary words. To allow for mistakes in drawing symbols and to allow the space needed for hand-drawn symbols to be judged accurately, you may wish to draw symbols on a separate sheet of paper and paste them into the text.

Hyphenation

Lines that contain large gaps between words or letters are difficult to read, so words should be broken between lines where necessary to avoid creating lines of this sort. Use a single h7yphen at the end of the line rather than a line under the last letter of the line. Dividing words in the proper places is important for readability. Use your dictionary.

Page Numbers

Pages should be numbered to ensure that they remain in the correct order. Numbers hould be placed outside th type area, at least 1/2 inch (13mm) from the edge of the type area.

CORRECTIONS

Corrected passages or pages introduced into the article must be indistinguishable from the rest of the article in type style, size, and darkness. Words must not be squeezed into the text, an dno unusual gaps should be left. The easiest and most effective method of introducing corrections is to correct the computer file and reprint the pertinent pages. If necessary, smaller segments of text may be reprinted and pasted into the article; paste up of corrections may be left to the publisher. Please check all reprinted pages or passages to ensure that they begin and end at the same point in the text as the elements they replace.

† Notes should not be included with the literature references. Footnotes should be set in type two

points smaller than the text type, but endnotes hould be set in ordinary text type.

FIGURES

To introduce uniformity of appearance to the volume, Plenum Press will scale all illustrations to be included in the book, produce sized prints, and mount the sized prints in the camera copy. Therefore, the original artwork or full-size photographic prints made from the original artwork should be supplied rather than reduced-size prints and should be supplied separately rather than being mounted in the article. Illustrations should be identified by author and figure number and the orientation should be indicated if it is not obvious. The illustrations will be scaled so that the capital letters of the bulk of the labels will measure about 3/32 inches (2.5 mm) in height. Please take this scaling into account when calculating how much space to leave for each figure in the text.

Figure Captions

A brief explanatory caption should appear below each figure. The caption should be printed in the typeface used in the text but two points smaller than the text type. Captions of figures that will be three-quarters page width or wider after scaling should be printed page width, and captions of narrower figures should be printed three-quarters page width.

Figure Placement and Spacing

Figures must be integrated with the text and not gathered at the end of the article. As there is no proof state in the camera-ready process, failure to integrate the figures into the text will make it impossible for the publisher to ensure that figures appear in the proper locations in the article. The figure caption and an appropriate amount of space for each figure should be placed near where the figure is discussed but not necessarily immediately after the line or paragraph in which it is first mentioned. If placing a figure on the page where it is first mentioned would require the creation of a short page, the figure should be placed on the preceding or following page. See Figure 1.

Figure 1. Example of a short figure caption.

Figures should be numbered consecutively throughout the article (1, 2, 3, …). In the text, figures should be identified by number rather than by “above,” “below,” etc. See Figure 2.

Figure 2. This is an example of a very, very, very long figure caption which takes up two lines of space.

A two-line space should be left below the caption, unless the caption is at the bottom of a page. In calculating how much space to leave for a figure, add in the one-line space that will divide the figure and caption and the two-line space that will separate the top of the figure from whatever is above it. Of course, if the figure will be at the top of a page, the two-line space should be omitted. Place figures side by side only if they will fit, when scaled, with an inch (25 mm) of space between them, and do not run text next to a figure. Large figures may be placed sideways (landscape orientation) on a page. All elements on a page containing a landscapeoriented material must be oriented int he same manner, and no text should be placed on such a page.

Line Drawings

Line drawings supplied should be original ink drawings, high-quality laser prints produced using appropriate software, or full-size photographic prints made from such artwork (see Table 2). If possible, drawings and laser prints should be prepared to be reduced by about 50% during the scaling process. Note that in labeling figures it is important to avoid making subscripts and superscripts too much smaller than the main characters; if the subscripts and superscripts are too small in comparison with the main type,, they will not be readable when the figures are reduced. If the original art cannot be obtained when reproducing a previously-published figure, the appropriate page from the publication involved or a photographic print made from the page may be supplied instead. Please do not supply xerographic copies.

Table 2. Suitability of artwork for reproduction.^{1}

Line drawings  Photographs to be reproduced as halftones
Good • Original ink drawings  • Continuous-tone photographic prints
• Crisp laser prints  • Continuous-tone negatives

• Photographic prints made from clear

originals

• Pages taken from publications

Bad  • Xerographic copies  • Xerographic copies
• Anything covered with transparent tape • Screened prints
• Drawings containing blurred or faded  • Pages taken from publications
lines, distorted symbols, etc.  (these are screened)

• Drawings containing faint or

damaged shading

1 Providing “good” artwork will help to speed the publication of the book.

Continuous-Tone Photographs

Photographs should be supplied either in the form of good-quality continuoustone photographic prints made form the original negatives or int eh form of the negatives themselves. Magnifications should be indicated by scale bars included the figures rather than stated in the captions.

Screened prints and xerographic copies will not reproduce well and are not acceptable.

Color Figures

It is technically feasible to reproduce figures in color, but it is expensive (perhaps $1800 per figure as of July, 1991). If a figure is to be reproduced in color, special arrangements must be made to cover the cost of color reproduction and specific instructions must be included indicating that color reproduction will be subsidized. Figures, originally in color, but to be reproduced in black and white in the book should be supplied in the form of black and white continuous-tone photographs made from the color originals. Such prints can be prepared by any photographic department or photography shop. Providing the figures in this form will give you the opportunity to see how well the figures wil reproduce in black and white, while obviating any misunderstanding regarding how the figures are to be reproduced. If you have a question about color reproduction, please get in touch with us at the earliest possible time.

TABLES

Table placement, spacing, and numbering should follow the guidelines given above for figures.

A table-width table title (caption), printed in type of the same size and style as the text type, should be placed above the table.

Large tables may be placed sideways on a page or may be photographically reduced to fit on the page. The table title should not be reduced.

Keeping the arrangement of tables simple will help to make the tables easy to read. In most cases, table-width rules at the top and bottom of the table body and below the column headings, if there are any, are all the lines necessary. Short straddle rules may be used to establish a hierarchy among column headings. Tables that are basically lists should have the same rules as other tables, viz., top and bottom rules, and rules below the column headings, if any.

It is a good practice to set tables (not, however, the titles) in type smaller than the text type. This sets the tables off from the text and helps to give a more uniform look to sets of tables that include large tables that would have to be set in smaller type anyway in order to fit on the page.

REFERENCES

1. R.J. Miller. “Book Title with the Initial Letter of Each Major Word Capitalized,” Publisher, City (1967).

2. A.B. Smith and C.D. Jones, Article or chapter title with only the first work having an initial capital, in: “Book Title,” W.F. White, ed., Publisher, City (1972).

3. C.D. Jones, M.P. Green, and S.A. Robinson, Article title, Journal Name Abbr. 37:468 (1974).